OpenStudy (anonymous):

The integral of ((4sin*squared*x cos*squared*x divided by sin2xcos2x))dx help :3

ganeshie8 (ganeshie8):

know ur trig

ganeshie8 (ganeshie8):

simplify the integrand first, this is more of a trig problem... not calculus :)

OpenStudy (anonymous):

i wouldn't post this if im good at my trig :D

ganeshie8 (ganeshie8):

il give a hint maybe : \(\large 2\sin (x) \cos( x) = \sin(2x)\)

ganeshie8 (ganeshie8):

does that help ? :)

OpenStudy (anonymous):

im trying :)

ganeshie8 (ganeshie8):

okie, take ur time :)

ganeshie8 (ganeshie8):

\(\large \int \frac{4\sin^2x \cos^2x}{\sin(2x) \cos(2x)} ~dx\)

ganeshie8 (ganeshie8):

the integrand is like that right ?

OpenStudy (anonymous):

yup

ganeshie8 (ganeshie8):

okay, take the numerator : \(\large 4\sin^2x \cos^2x\) \(\large (2\sin x \cos x)^2\) appeal to the earlier identity(hint) : \(\large \left(\sin (2x)\right)^2\)

ganeshie8 (ganeshie8):

see if that looks okay...

ganeshie8 (ganeshie8):

then, the integrand becomes : \(\large \frac{\left(\sin(2x)\right)^2}{\sin(2x) \cos(2x)} \)

OpenStudy (anonymous):

so the answer will be -1/2 ln *absolute value of* cos2x + C???

ganeshie8 (ganeshie8):

Looks you nailed it !!

ganeshie8 (ganeshie8):

Yup, thats the correct answer :)

OpenStudy (anonymous):

thank you :D

ganeshie8 (ganeshie8):

np... u wlc :)